In a science test, the mean score is 42 and the standard deviation is 5. Assuming the scores are normally distributed, what percent of the score is greater than 47


1
Expert's answer
2022-03-28T16:05:40-0400

"\mu=42,\sigma=5"

"P(x>47)=1-P(x<47)"

"=1-P(Z<\frac{47-42}{5})"

"=1-P(Z<1.0)"

"=1-0.8413"

"=0.1587"